OpenStudy (anonymous):

Given that q = −1.60 µC and d = 20.0 cm, find the electric potential energy of the three charges system

OpenStudy (ookawaiioo):

We know that Force of a charge (Fq) = q1*q2/r^2 Distance is 20, then r = 10. q1 & q2 = −1.60 µC PE = work = Force* distance So (Fq) = q1*q2/r <----- use this equation. I skipped a step because Im lazy.